# Kaon-induced production of strange hidden-charm molecular pentaquarks $P^{\Lambda}_{\psi s}$ from proton

### v1

@@ Abstract
In this work, we investigate the kaon-induced production of strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}$ states, including the $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4338)$ observed by the LHCb collaboration. Coupled-channel interactions involving the channels $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$ are studied using effective Lagrangians to describe these interactions. The potential kernels are constructed within the one-boson-exchange model and implemented into the quasipotential Bethe-Salpeter equation to calculate the scattering amplitudes. From these amplitudes, the partial-wave cross sections for kaon-induced production are predicted alongside the poles corresponding to strange hidden-charm molecular pentaquarks. The analysis reveals complex structures in the cross sections due to molecular states. A structure near the $\Xi_{c}\bar{D}$ threshold is predicted in the $\Lambda_{c}\bar{D}_{s}$ channel, corresponding to $P^{\Lambda}_{\psi s}(4338)$. Both molecular states near the $\Xi_{c}\bar{D}^{*}$ threshold with $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$, associated with $P^{\Lambda}_{\psi s}(4459)$, produce peaks in the $\Xi^{\prime}_{c}\bar{D}$ channel, which vary significantly with the cutoff. Additionally, a molecular state with $J^{P}=3/2^{-}$ near the $\Xi_{c}^{*}\bar{D}$ threshold gives rise to a peak in the $\Xi_{c}\bar{D}^{*}$ channel. In the higher energy region, from 4350 to 4550 MeV, although the cross sections in the $\Lambda_{c}\bar{D}_{s}^{(*)}$ channels are substantial, no distinct structures corresponding to strange hidden-charm pentaquarks are identified, apart from some threshold-related cusps. These results provide valuable theoretical insights and serve as a foundation for future experimental studies of strange hidden-charm pentaquarks using kaon beams.
I INTRODUCTION
@@
In 2020, the LHCb Collaboration reported the observation of the strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}(4459)$ in the $\Xi_{b}^{-}\to J/\psi\Lambda K^{-}$ decay Aaij:2020gdg . This state has a measured mass approximately 19 MeV below the $\Xi_{c}\bar{D}^{*}$ threshold and a width of 17 MeV Aaij:2020gdg , characteristics consistent with its interpretation as a molecular state. Numerous studies have investigated the molecular state nature of $P^{\Lambda}_{\psi s}(4459)$. For instance, our previous work suggests that the structure of $P^{\Lambda}_{\psi s}(4459)$ primarily arises from the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=1/2^{-}$, while the possible contribution of the $\Xi_{c}\bar{D}^{*}$ state with $J^{P}=3/2^{-}$ cannot be excluded Zhu:2021lhd ; Zhu:2022wpi . Similarly, QCD sum rule analyses in Ref. Chen:2020uif support the interpretation of $P^{\Lambda}_{\psi s}(4459)$ as a $\Xi_{c}\bar{D}^{*}$ molecular state with $J^{P}=1/2^{-}$ or $3/2^{-}$. Ref. Wang:2021itn interprets $P^{\Lambda}_{\psi s}(4459)$ as either a $\Xi^{\prime}_{c}\bar{D}$ molecular state with quantum numbers $I(J^{P})=0(3/2^{-})$ or a $\Xi_{c}\bar{D}^{*}$ molecular state with $I(J^{P})=0(3/2^{-})$. Moreover, studies within the framework of heavy quark spin symmetry (HQSS), such as those in Ref. Peng:2020hql , favor the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=3/2^{-}$ as the most likely candidate for $P^{\Lambda}_{\psi s}(4459)$, ruling out the $J^{P}=1/2^{-}$ possibility.
@@
Although hidden-charm pentaquarks, including those with strangeness, have been observed at LHCb, further confirmation and a deeper understanding of these states require analyzing their production mechanisms across different reaction channels. Even before the observation of the $P^{N}_{\psi}$ state at LHCb, the production of these states via proton-antiproton collisions and photoproduction was proposed Wu:2010jy ; Huang:2013mua ; He:2015cea . Following the LHCb observations, many studies were proposed to investigate the production of hidden-charm pentaquarks, $P^{\Lambda}_{\psi}$, via pion-induced reactions or photoproduction Lu:2015fva ; Huang:2016tcr ; Kim:2016cxr ; Cao:2019gqo ; Wang:2019dsi ; Wang:2019krd ; Xie:2020niw ; Shi:2022ipx . Additionally, the pion-induced production of strange hidden-charm pentaquarks was studied even before the LHCb observation Wang:2015xwa . In Refs. Clymton:2022qlr ; Clymton:2021thh ; Paryev:2023icm ; Paryev:2022zdx , the authors employed two different methods, the effective Lagrangian approach and the Regge method, to study the production of the exotic hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ in the $\bar{K}p\rightarrow J/\psi\Lambda$ reaction. In Ref. Cheng:2021gca , the authors explored the possibility of investigating the strange hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ through photon-induced reactions on a proton target in the $\gamma p\rightarrow K^{+}P^{\Lambda}_{\psi s}(4459)$ channel, using an effective Lagrangian approach.
@@
This article is organized as follows. After the introduction, Sec. II outlines the Lagrangians used to construct the potentials for coupled-channel interactions, the qBSE approach, and the methodology for calculating the scattering cross sections. In Sec. III, we present the results for the cross sections of kaon-induced production. Finally, Sec. IV provides a summary of the entire work.
II Formalism
@@
II.1 Relevant lagrangians and potential kernels
In the present work, we consider molecular pentaquarks produced from the interaction of charmed baryons and anticharmed mesons, including the channels $\Xi_{c}^{*}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}^{*}$, $\Xi^{*}_{c}\bar{D}$, $\Xi_{c}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}$, $\Lambda_{c}\bar{D}_{s}$, $\Xi_{c}\bar{D}$, and $\Lambda_{c}\bar{D}_{s}$, with light meson exchange within the one-boson-exchange model. We also consider the channel involving a hidden-charm meson, $\Lambda J/\psi$, for which the interaction is neglected. Therefore, we require vertices for the coupling between the constituent particles and the exchanged meson. The relevant Lagrangians for these vertices are given as follows Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Casalbuoni:1996pg ; Zhu:2021lhd ,
@@ II.1 Relevant lagrangians and potential kernels
$\mathcal{L}_{\bar{K}\bar{K}{V}}$ $=ig_{\bar{K}\bar{K}V}\bar{K}^{{\dagger}}{V}^{\mu}\overleftrightarrow{\partial}_{\mu}\bar{K},$
@@
[…]
@@
When calculating the coupling interaction between $\bar{K}N$ and other channels, in addition to the Lagrangians mentioned above, we also need to use the following Lagrangians Wang:2016fhj ; Okubo:1975sc ; Sibirtsev:2000aw ; Liu:2001ce ; Dong:2009tg ; Lu:2016nnt ,
@@
Following the method in Ref. He:2019rva , we input the vertices $\Gamma$ and propagators $P$ directly into the code. The explicit forms of the potential can be written using the Lagrangians and flavor wave functions as follows,
@@ II.2 The qBSE approach
where the summation is restricted to nonnegative helicities $\lambda^{\prime\prime}$. The propagator $G_{0}({\rm p}^{\prime\prime})$ is simplified from the original 4-dimensional form under the quasipotential approximation and takes the form $G_{0}({\rm p}^{\prime\prime})=\delta^{+}(p^{\prime\prime 2}_{h}-m_{h}^{2})/(p^{\prime\prime 2}_{l}-m_{l}^{2})$, where $p^{\prime\prime}_{h,l}$ and $m_{h,l}$ represent the momenta and masses of the heavy and light constituent particles, respectively. The partial wave potential is expressed as:
@@
where $\eta=PP_{1}P_{2}(-1)^{J-J_{1}-J_{2}}$, with $P$ and $J$ denoting the parity and spin of the system, as well as those of constituent particles 1 and 2. The relative momenta for the initial and final states are specified as ${\bm{p}}=(0,0,{\rm p})$ and ${\bm{p}}^{\prime}=({\rm p}^{\prime}\sin\theta,0,{\rm p}^{\prime}\cos\theta)$, respectively. The term $d^{J}_{\lambda\lambda^{\prime}}(\theta)$ represents the Wigner d-matrix, and the amplitude integration is given by,
[…]
@@
$G_{j=0}$ $=$ $-\frac{i{\rm p}^{\prime\prime}_{o}}{32\pi^{2}W}+\sum_{j}\left[\frac{w({\rm p}_{j})}{(2\pi)^{3}}\frac{{\rm p}^{\prime\prime 2}_{o}}{2W{({\rm p}^{\prime\prime 2}_{j}-{\rm p}^{\prime\prime 2}_{o})}}\right].$ (30)
The on-shell momentum is given by ${\rm p}^{\prime\prime}_{o}=\lambda^{1/2}(W,M_{1},M_{2})/2W$, where $\lambda(x,y,z)=[x^{2}-(y+z)^{2}][x^{2}-(y-z)^{2}]$ and $W$ represents the total energy of the two-constituent system. To regularize the propagator, an exponential form factor is introduced as $G_{0}({\rm p}^{\prime\prime})\to G_{0}({\rm p}^{\prime\prime})\left[e^{-(p^{\prime\prime 2}_{l}-m_{l}^{2})^{2}/\Lambda_{r}^{4}}\right]^{2}$, where $\Lambda_{r}$ denotes the cutoff He:2015mja . In our framework, the cutoffs applied in the form factors, including those in the propagator and exchange interactions, are treated as free parameters, with $\Lambda_{e}=\Lambda_{r}=\Lambda$. We adopt a cutoff value of $\Lambda=1.04$ GeV, which was successfully used to fit the LHCb data in our previous woork Zhu:2022wpi , allowing for slight variations of 0.1 GeV to estimate uncertainties.
The poles of the nine-channel scattering amplitudes for different spin-parity combinations $J^{P}$ are located by varying $z$ in the complex energy plane to satisfy the condition $|1-V(z)G(z)|=0$. For a calculation involving $N$ channels, there are $2^{N}$ Riemann sheets associated with unitarity. In this study, we adopt the method outlined in Ref. Roca:2005nm to identify the poles. Specifically, the default propagator $G(z)$ is used for the energy region below the threshold, while $G(z)+{i{\rm p}^{\prime\prime}_{o}}/{16\pi^{2}z}$ is applied for the region above the threshold, corresponding to the first and second Riemann sheets, respectively.
@@
III Numerical Result
III.1 Molecular states without coupled-channel effects

### v2

@@ Abstract
In this work, we investigate the kaon-induced production of strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}$ states, including the $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4338)$ observed by the LHCb collaboration. Coupled-channel interactions involving the channels $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$ are studied using effective Lagrangians to describe these interactions. The potential kernels are constructed within the one-boson-exchange model and implemented into the quasipotential Bethe-Salpeter equation to calculate the scattering amplitudes. From these amplitudes, the partial-wave cross sections for kaon-induced production are predicted alongside the poles corresponding to strange hidden-charm molecular pentaquarks. The analysis reveals complex structures in the cross sections due to molecular states. A structure near the $\Xi_{c}\bar{D}$ threshold is predicted in the $\Lambda_{c}\bar{D}_{s}$ channel, corresponding to $P^{\Lambda}_{\psi s}(4338)$. Both molecular states near the $\Xi_{c}\bar{D}^{*}$ threshold with $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$, associated with $P^{\Lambda}_{\psi s}(4459)$, produce peaks in the $\Xi^{\prime}_{c}\bar{D}$ channel, which vary significantly with the cutoff. Additionally, a molecular state with $J^{P}=3/2^{-}$ near the $\Xi_{c}^{*}\bar{D}$ threshold gives rise to a peak in the $\Xi_{c}\bar{D}^{*}$ channel. In the higher energy region, from 4350 to 4550 MeV, although the cross sections in the $\Lambda_{c}\bar{D}_{s}^{(*)}$ channels are substantial, no distinct structures corresponding to strange hidden-charm pentaquarks are identified, apart from some threshold-related cusps. These results provide valuable theoretical insights and serve as a foundation for future experimental studies of strange hidden-charm pentaquarks using kaon beams at J-PARC and JLab.
journal: Eur. Phys. J. C
1 INTRODUCTION
@@
In 2020, the LHCb Collaboration reported the observation of the strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}(4459)$ in the $\Xi_{b}^{-}\to J/\psi\Lambda K^{-}$ decay Aaij:2020gdg . This state has a measured mass approximately 19 MeV below the $\Xi_{c}\bar{D}^{*}$ threshold and a width of 17 MeV Aaij:2020gdg , characteristics consistent with its interpretation as a molecular state. Numerous studies have investigated the molecular state nature of $P^{\Lambda}_{\psi s}(4459)$. For instance, our previous work suggests that the structure of $P^{\Lambda}_{\psi s}(4459)$ primarily arises from the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=1/2^{-}$, while the possible contribution of the $\Xi_{c}\bar{D}^{*}$ state with $J^{P}=3/2^{-}$ cannot be excluded Zhu:2021lhd ; Zhu:2022wpi . Similarly, QCD sum rule analyses in Ref. Chen:2020uif support the interpretation of $P^{\Lambda}_{\psi s}(4459)$ as a $\Xi_{c}\bar{D}^{*}$ molecular state with $J^{P}=1/2^{-}$ or $3/2^{-}$. Ref. Wang:2021itn interprets $P^{\Lambda}_{\psi s}(4459)$ as either a $\Xi^{\prime}_{c}\bar{D}$ molecular state with quantum numbers $I(J^{P})=0(3/2^{-})$ or a $\Xi_{c}\bar{D}^{*}$ molecular state with $I(J^{P})=0(3/2^{-})$. Moreover, studies within the framework of heavy quark spin symmetry (HQSS), such as those in Ref. Peng:2020hql , favor the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=3/2^{-}$ as the most likely candidate for $P^{\Lambda}_{\psi s}(4459)$, ruling out the $J^{P}=1/2^{-}$ possibility. A recent study suggests that the strong coupling between $\bar{D}_{s}\Lambda_{c}$ and $\bar{D}\Xi_{c}$, as well as between $\bar{D}s^{*}\Lambda_{c}$ and $\bar{D}^{*}\Xi_{c}$, leads to the association of $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$ with states predominantly coupling to $\bar{D}\Xi^{\prime}$ and $\bar{D}^{*}\Xi_{c}$, respectively Feijoo:2022rxf .
@@
Although hidden-charm pentaquarks, including those with strangeness, have been observed at LHCb, further confirmation and a deeper understanding of these states require analyzing their production mechanisms across different reaction channels. Even before the observation of the $P^{N}_{\psi}$ state at LHCb, the production of these states via proton-antiproton collisions and photoproduction was proposed Wu:2010jy ; He:2015cea ; Huang:2013mua . Following the LHCb observations, many studies were proposed to investigate the production of hidden-charm pentaquarks, $P^{\Lambda}_{\psi}$, via pion-induced reactions or photoproduction Lu:2015fva ; Huang:2016tcr ; Kim:2016cxr ; Cao:2019gqo ; Wang:2019krd ; Xie:2020niw ; Shi:2022ipx ; Wang:2019dsi . Additionally, the pion-induced production of strange hidden-charm pentaquarks was studied even before the LHCb observation Wang:2015xwa . In Refs. Clymton:2021thh ; Paryev:2023icm ; Paryev:2022zdx ; Clymton:2022qlr , the authors employed two different methods, the effective Lagrangian approach and the Regge method, to study the production of the exotic hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ in the $\bar{K}p\rightarrow J/\psi\Lambda$ reaction. In Ref. Cheng:2021gca , the authors explored the possibility of investigating the strange hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ through photon-induced reactions on a proton target in the $\gamma p\rightarrow K^{+}P^{\Lambda}_{\psi s}(4459)$ channel, using an effective Lagrangian approach.
@@
This article is organized as follows. After the introduction, Sec. 2 outlines the Lagrangians used to construct the potentials for coupled-channel interactions, the qBSE approach, and the methodology for calculating the scattering cross sections. In Sec. 3, we present the results for the cross sections of kaon-induced production. Finally, Sec. 4 provides a summary of the entire work.
2 Formalism
@@
2.1 Relevant lagrangians and potential kernels
In the present work, we consider molecular pentaquarks produced from the interaction of charmed baryons and anticharmed mesons, including the channels $\Xi_{c}^{*}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}^{*}$, $\Xi^{*}_{c}\bar{D}$, $\Xi_{c}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}$, $\Lambda_{c}\bar{D}_{s}$, $\Xi_{c}\bar{D}$, and $\Lambda_{c}\bar{D}_{s}$, with light meson exchange within the one-boson-exchange model. We also consider the channel involving a hidden-charm meson, $\Lambda J/\psi$, for which the interaction is neglected. Therefore, we require vertices for the coupling between the constituent particles and the exchanged meson. The relevant Lagrangians for these vertices are given as follows Zhu:2021lhd ; Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Casalbuoni:1996pg ,
@@ 2.1 Relevant lagrangians and potential kernels
Based on the aforementioned Lagrangians, we can derive the interaction mechanisms responsible for generating molecular strange hidden-charm pentaquarks. To investigate their kaon production dynamics, it is crucial to incorporate couplings between these exotic states and the $\bar{K}N$ channel. While the broad energy region covered here suggests potential contributions from multiple inelastic channels, our current focus remains on providing experimentally testable predictions. Notably, these inelastic channels primarily influence cross-section magnitudes rather than the fundamental production mechanisms. Given the current absence of experimental data for these exotic states, our theoretical predictions target order-of-magnitude estimates - a resolution level sufficient for guiding initial experimental searches. Moreover, the essential characteristics of cross-section behavior (excluding absolute normalization) remain relatively insensitive to these coupling details due to the large mass gap. We therefore employ a simplified one-boson exchange framework to model the crucial coupling between the $\bar{K}N$ channel and charmed-sector final states, maintaining theoretical consistency while preserving predictive capability.
@@
$\mathcal{L}_{\bar{K}\bar{K}{\rho}}$ $=ig_{\bar{K}\bar{K}\rho}\bar{K}^{{\dagger}}{\bm{\tau}}\cdot{\bm{\rho}}^{\mu}\overleftrightarrow{\partial}_{\mu}\bar{K},$
$\mathcal{L}_{\bar{K}\bar{K}{\omega}}$ $=ig_{\bar{K}\bar{K}\omega}\bar{K}^{{\dagger}}{\omega}^{\mu}\overleftrightarrow{\partial}_{\mu}\bar{K},$
@@
$\mathcal{L}_{NN\pi}$ $=-\frac{g_{NN\pi}}{m_{\pi}}\bar{N}\gamma^{5}\gamma^{\mu}{\bm{\tau}}\cdot\partial_{\mu}{\bm{\pi}}N,\quad$
$\mathcal{L}_{NN\rho}$ $=-g_{NN\rho}\bar{N}[\gamma^{\mu}-\frac{\kappa_{\rho}}{2m_{N}}\sigma^{\mu\nu}\partial_{\nu}]{\bm{\tau}}\cdot{\bm{\rho}_{\mu}}N,$
$\mathcal{L}_{NN\omega}$ $=-g_{NN\omega}\bar{N}[\gamma^{\mu}-\frac{\kappa_{\omega}}{2m_{N}}\sigma^{\mu\nu}\partial_{\nu}]{\omega_{\mu}}N,$
@@
When calculating the couplings between $\bar{K}N$ and other channels, in addition to the Lagrangians mentioned above, we also need to use the following Lagrangians Wang:2016fhj ; Okubo:1975sc ; Sibirtsev:2000aw ; Liu:2001ce ; Dong:2009tg ; Lu:2016nnt ,
@@
In the current framework, the inclusion of nine reaction channels naturally leads to 81 pairwise interactions requiring calculation. Explicitly enumerating all these interaction terms would not only prove exceptionally tedious but also introduce significant error-prone risks in formulation. Following the method in Ref. He:2019rva , we input the vertices $\Gamma$ and propagators $P$ directly into the code. The explicit forms of the potential can be written using the Lagrangians and flavor wave functions as follows,
@@ 2.2 The qBSE approach
where the summation is restricted to nonnegative helicities $\lambda^{\prime\prime}$. The propagator $G_{0}({\rm p}^{\prime\prime})$ is simplified from its original four-dimensional form under the quasipotential approximation and takes the form
$G_{0}$ $=\frac{\delta^{+}(p^{\prime\prime~{}2}_{h}-m_{h}^{2})}{p^{\prime\prime 2}_{l}-m_{l}^{2}}$
$=\frac{\delta^{+}(p^{\prime\prime 0}_{h}-E_{h}({\rm p}^{\prime\prime}))}{2E_{h}({\rm p}^{\prime\prime})[(W-E_{h}({\rm p}^{\prime\prime}))^{2}-E_{l}^{2}({\rm p}^{\prime\prime})]}.$ (27)
As required by the spectator approximation adopted in this work, the heavier particle (denoted as $h$) in a given channel is placed on shell Gross:1999pd , which satisfies $p^{\prime\prime 0}_{h}=E_{h}({\rm p}^{\prime\prime})=\sqrt{m_{h}^{2}+{\rm p}^{\prime\prime 2}}$. The energy of the lighter particle (denoted as $l$) is then given by $p^{\prime\prime 0}_{l}=W-E_{h}({\rm p}^{\prime\prime})$. Here and hereafter, the value of the momentum in center-of-mass frame is defined as ${\rm p}=|{\bm{p}}|$. The partial wave potential is expressed as:
@@
where $\eta=PP_{1}P_{2}(-1)^{J-J_{1}-J_{2}}$, with $P$ and $J$ denoting the parity and spin of the system, as well as those of constituent particles 1 and 2. The relative momenta for the initial and final states are specified as ${\bm{p}}=(0,0,{\rm p})$ and ${\bm{p}}^{\prime}=({\rm p}^{\prime}\sin\theta,0,{\rm p}^{\prime}\cos\theta)$, respectively. The term $d^{J}_{\lambda\lambda^{\prime}}(\theta)$ represents the Wigner d-matrix.
@@
$G_{j=0}$ $=-\frac{i{\rm p}^{\prime\prime}_{\rm o}}{32\pi^{2}W}+\sum_{j}\left[\frac{w({\rm p}_{j})}{(2\pi)^{3}}\frac{{\rm p}^{\prime\prime 2}_{\rm o}}{2W{({\rm p}^{\prime\prime 2}_{j}-{\rm p}^{\prime\prime 2}_{\rm o})}}\right].$ (30)
The ${\rm p}^{\prime\prime}_{\rm o}=\lambda^{1/2}(W,M_{1},M_{2})/2W$ is on-shell momentum with $\lambda(x,y,z)=[x^{2}-(y+z)^{2}][x^{2}-(y-z)^{2}]$ and $W$ represents the total energy of the two-constituent system. To regularize the propagator, an exponential form factor is introduced as $G_{0}({\rm p}^{\prime\prime})\to G_{0}({\rm p}^{\prime\prime})\left[e^{-(p^{\prime\prime 2}_{l}-m_{l}^{2})^{2}/\Lambda_{r}^{4}}\right]^{2}$, where $\Lambda_{r}$ denotes the cutoff He:2015mja . In our framework, the cutoffs applied in the form factors, including those in the propagator and exchange interactions, are treated as free parameters, with $\Lambda_{e}=\Lambda_{r}=\Lambda$. We adopt a cutoff value of $\Lambda=1.04$ GeV, which was successfully used to fit the LHCb data in our previous work Zhu:2022wpi , allowing for slight variations of 0.1 GeV to estimate uncertainties.
The poles of the nine-channel scattering amplitudes for different spin-parity combinations $J^{P}$ are located by varying $z$ in the complex energy plane to satisfy the condition $|1-V(z)G(z)|=0$. For a calculation involving $N$ channels, there are $2^{N}$ Riemann sheets associated with unitarity. In this study, we adopt the method outlined in Ref. Roca:2005nm to identify the poles. Specifically, the default propagator $G(z)$ is used for the energy region below the threshold, while $G(z)+i{\rm p}^{\prime\prime}_{o}/(16\pi^{2}z)$ is adopted for the region above the threshold, corresponding to the first and second Riemann sheets, respectively. This treatment does not affect the physical results on the real energy axis and yields pole positions and half-widths that are closer to those of the corresponding Breit-Wigner forms.
@@
3 Numerical Result
3.1 Molecular states without coupled-channel effects